OpenStudy (anonymous):

Two vessels contain milk mixed with water in the ratio 7:3 and 6:5 . The quantity of milk solution (in litres) from the first vessel to be mixed with 33 litres from the second so that the ratio of milk and water becomes 46:27 is?

OpenStudy (anonymous):

@ajprincess @apoorvk @amistre64

OpenStudy (anonymous):

@satellite73

OpenStudy (phi):

ratio of 7:3 mean 7/10 of the solution is milk we have 0.7 x + 6*33/11 = 46y/73 also x+33= y solve for x

OpenStudy (anonymous):

40?

OpenStudy (phi):

that is what I got

OpenStudy (anonymous):

The options 35 40 45 50

OpenStudy (anonymous):

40 :P

OpenStudy (anonymous):

@phi since there r two unknown variable how will we solve it

OpenStudy (phi):

you have 2 equations. the 2nd is x+33= y (the amount of the 1st solution plus 33 liters from the second solution add up to a total of y)

OpenStudy (anonymous):

@phi how x+33= y???

OpenStudy (phi):

The quantity of milk solution (in litres) from the first vessel (CALL THIS X) to be mixed with 33 litres from the second so X+33= Y (Y is the 2nd unknown)